# SportR: A Benchmark for Multimodal Large Language Model Reasoning in Sports

### v2

@@ Abstract
Artificial Intelligence brings powerful new tools to sports, from automated officiating to tactical analysis, but these applications all depend on a core reasoning capability. Deeply understanding sports requires an intricate blend of fine-grained visual perception and rule-based reasoning—a challenge that pushes the limits of current multimodal models. To succeed, models must master three critical capabilities: perceiving nuanced visual details, applying abstract sport rule knowledge, and grounding that knowledge in specific visual evidence. Current sports benchmarks either cover single sports or lack the detailed reasoning chains and precise visual grounding needed to robustly evaluate these core capabilities in a multi-sport context. To address this gap, we introduce SportR, the first multi-sports large-scale benchmark designed to train and evaluate MLLMs on the fundamental reasoning required for sports intelligence. Our benchmark provides a dataset of 5,017 images and 2,101 videos. To enable granular evaluation, we structure our benchmark around a progressive hierarchy of question-answer (QA) pairs designed to probe reasoning at increasing depths—from simple infraction identification to complex penalty prediction. For the most advanced tasks requiring multi-step reasoning, such as determining penalties or explaining tactics, we provide 7,118 high-quality, human-authored Chain-of-Thought (CoT) annotations. In addition, our benchmark incorporates both image and video modalities and provides manual bounding box annotations to test visual grounding in the image part directly. Extensive experiments demonstrate the profound difficulty of our benchmark. State-of-the-art baseline models perform poorly on our most challenging tasks. While training on our data via Supervised Fine-Tuning and Reinforcement Learning improves these scores, they remain relatively low, highlighting a significant gap in current model capabilities. SportR presents a new challenge for the community, providing a critical resource to drive future research in multimodal sports reasoning.
@@ 1 Introduction
To address these critical gaps, we introduce SportR, a new large-scale, multimodal benchmark designed to train and evaluate fine-grained reasoning for fundamental sports understanding. We focus on a diverse set of five globally popular ball and racket sports – basketball, soccer, table tennis, badminton, and American football – to provide a rich testbed for generalization. Our benchmark is based on a corpus of 5,017 images (from all five sports) and 2,101 videos (from four sports, excluding badminton), encompassing 50 distinct foul types and 12 fundamental tactics. The cornerstone of our work is a collection of 7,118 high-quality, fully human-annotated CoT. We also derive over 20,000 structured question-answer pairs.
@@ 2.2 Multimodal Sports QA Benchmarks
Question answering (QA) has become the standard way for evaluating the reasoning abilities of MLLMs across different domains Chen et al. (2025a); Liu et al. ; Yue et al. (2024); Wu et al. (2023); Shao et al. (2024a). In the sports domain, early multimodal QA benchmarks primarily focused on assessing models’ ability to recognize actions and answer descriptive or temporal questions. For instance, Sports-QA (Li et al., 2024c) leveraged existing action recognition datasets to create a large-scale VQA benchmark, but its questions do not center on the complex, rule-based scenarios that define competitive play.
More recent benchmarks have begun to address this gap by focusing on deeper forms of reasoning. Notably, SPORTU (Xia et al., 2024b) introduced a comprehensive evaluation benchmark with a multi-level difficulty design, explicitly including “hard” questions that require understanding game rules and tactics. However, while SPORTU provides explanations for evaluation, the limitations are that there are not enough of this type of question, and these rationales were not designed as detailed and fine-grained, human-annotated reasoning processes suitable for training models to perform explicit, step-by-step reasoning. Furthermore, it lacks precise annotations for visual grounding, making it difficult to assess whether a model’s judgment is based on specific visual evidence, such as the subtle point of contact defining a foul.
@@ 3 SportR Benchmark
SportsImage – a dataset of 5,017 images designed to test a model’s ability to connect abstract rules to precise evidence in static scenes. Each image is paired with a CoT rationale focused on a specific foul or tactical play. This component’s progressive QA suite culminates in a unique visual grounding task, where the model is required to output the exact bounding box coordinates of the rule infraction.
SportsVideo – a dataset of 2,101 video clips created to extend the reasoning challenge to the temporal domain. This component addresses scenarios where understanding context is only possible by observing motion over time, such as a “traveling” violation in basketball. Each video is also annotated with a comprehensive CoT, from which a similar suite of progressive QA pairs. This part aims to provide a more thorough evaluation benchmark in sports understanding.
Overall, SportR contains 7,118 unique, human-authored CoT rationales, from which we derive over 20,000 structured question-answer pairs. This benchmark fills a critical gap by providing the resources to train models not just to answer what happened, but to reason about why it happened based on explicit, grounded visual evidence.
Our design prioritizes not only assessing a model’s reasoning but also verifying that this reasoning is matched to specific visual evidence. To this end, we introduce a novel explicit grounding task, where models must output the precise bounding box coordinates of a rule infraction. We focus this coordinate-based evaluation on SportsImage because fouls are often defined by discrete, localized events—such as a specific point of physical contact—that can be unambiguously annotated and evaluated in a static frame. This provides a clean and tractable setting to evaluate a skill that, as we later demonstrate, poses a significant challenge even for state-of-the-art models. Establishing a reliable baseline for this explicit form of grounding is a critical first step before tackling the greater spatio-temporal complexity of localization in video, which remains an important direction for future work. Examples of our SportsImage and SportsVideo can be found in Appendix C
@@ 3.1 Quality Control
A key principle of our data curation was to focus on fundamental and widely understood fouls and tactics—such as a basketball blocking foul or a standard post move—rather than highly specialized or controversial professional-level edge cases. This focus ensures that the scenarios, while challenging for AI systems, are grounded in the established, practical expertise of our annotation team. This allows for the creation of clear, consistent, and authoritative rationales, minimizing ambiguity and subjectivity.
The process began with a training phase where all annotators collaboratively developed and standardized the guidelines for writing CoT rationales and annotating grounding boxes. Each member then worked on small batches of examples, with the group reviewing the outputs to ensure a unified standard of quality and detail before commencing the full-scale annotation.
@@ 3.2 SportsImage: Multimodal Image QA
SportsImage is the first component of our benchmark, designed to rigorously evaluate an MLLM’s ability to perform fine-grained reasoning and rule-based visual grounding in static scenes. This part covers five sports: basketball, soccer, table tennis, badminton, and American football. It comprises 5,017 images, each capturing a critical moment of gameplay across multiple sports.
Each image is paired with a comprehensive, human-authored CoT rationale. This rationale serves as the ground truth to generate a progressive QA suite of up to seven distinct questions, resulting in over 16,000 structured question-answer pairs for this component alone. These questions are structured to probe a model’s reasoning at increasing levels of depth, covering: (Q1) infraction identification, (Q2) foul classification, (Q3) penalty prediction, (Q4) free-form explanation, (Q6) offensive tactic identification, (Q7) defensive tactic identification, and culminating in (Q5) an explicit visual grounding task requiring the model to output precise bounding box coordinates.
@@ 3.3 SportsVideo: Multimodal Video QA
The dataset consists of 2,101 video clips, from which we derive over 6,000 structured question-answer pairs. Like its image-based counterpart, each video is annotated with a comprehensive, human-authored CoT rationale that explains the core foul or tactic. This CoT then serves as the ground truth for a progressive QA suite designed to evaluate spatio-temporal reasoning.
@@ 4.1.1 Model Training on SportR
Beyond zero-shot baseline evaluations, we fine-tuned an open-source model to validate SportR as an effective training resource. We employed a two-stage training process on the Qwen-2.5VL-7B model: an initial Supervised Fine-Tuning (SFT) phase followed by Reinforcement Learning (RL) using the Group Relative Policy Optimization (GRPO) algorithm (Shao et al., 2024b). The training was conducted exclusively on our SportsImage component to demonstrate its value for teaching fine-grained visual reasoning. Full details on our data preparation, SFT/RL methodology, and reward function design are deferred to Appendix A.
@@ 4.2 Evaluation
We employ two primary metrics: Intersection over Union (IoU) (Everingham et al., 2010) for the visual grounding task (Q5) and an LLM-as-Judge framework for all other text-based QA tasks (detaile prompts are in Appendix B. To mitigate the known issue of self-preference bias where models may unfairly favor their own outputs (Zheng et al., 2023), we use three proprietary models as judges: GPT-5 (a20, 2025), Gemini 2.5 Pro (Comanici et al., 2025), and Claude 4.0 Sonnet (Anthropic, 2025) and report the average score. For each generated response, we report the scores from all three judge models as well as the average score to provide a comprehensive and robust assessment of performance.
@@ 5 Result
Our experiments are designed to demonstrate both the profound challenge posed by our benchmark and its utility as a training resource. We present the results for the SportsImage and SportsVideo components separately to provide analysis of the model’s capabilities in both static and video settings.
@@ 5.1 Performance on SportsImage
The zero-shot results underscore the profound difficulty of our benchmark. Even the most powerful proprietary model, GPT-5, struggles to achieve high accuracy on the core reasoning tasks. Across all baselines, performance is particularly low on tasks requiring deep, specialized knowledge, such as Foul Classification (Q2) and, most notably, the Explicit Visual Grounding task (Q5), where IoU scores are consistently below 7%. This confirms that existing models lack the fine-grained perception and rule-based reasoning abilities that SportR is designed to measure.
@@ 5.2 Performance on SportsVideo
Table 2: Performance comparison across models (%) on SportsVideo. Q8: Video Infraction identification; Q9: Video Foul Classification; Q10: Video Penalty Prediction; Q11: Video Free-form Explanation. Q12: Video Offensive Tactic Identification; Q13: Video Defensive Tactic Identification.
@@ 6 Conclusion
In this paper, we introduced SportR, a large-scale benchmark designed to train and evaluate the fine-grained, rule-based visual reasoning of Multimodal Large Language Models. Built upon a foundation of over 7,000 fully human-authored Chain-of-Thought rationales, we aim to provide a benchmark that can be used both for training and evaluation. By integrating tasks across both images and videos, our benchmark provides a holistic assessment of a model’s ability in sports understanding. Our experiments demonstrate that while our dataset is an effective training resource, it presents a profound challenge to current models. For example, even after fine-tuning and reinforcement learning, performance on the explicit grounding task remains modest, underscoring the difficulty of connecting abstract rules to precise visual evidence. We hope SportR will serve as a critical tool for the community, inspiring advancements in MLLMs and contributing to more robust and reliable real-world sports understanding.
Appendix A Detailed Model Training Methodology
@@
A.1 Model Training on SportR
@@
A.2 Training Data Preparation

### v3

@@ Abstract
Artificial Intelligence brings powerful new tools to sports, from automated officiating to tactical analysis, but these applications all depend on a core reasoning capability. Deeply understanding sports requires an intricate blend of fine-grained visual perception and rule-based reasoning—a challenge that pushes the limits of current multimodal models. To succeed, models must master three critical capabilities: perceiving nuanced visual details, applying abstract sport rule knowledge, and grounding that knowledge in specific visual evidence. Current sports benchmarks either cover single sports or lack the detailed reasoning chains and precise visual grounding needed to robustly evaluate these core capabilities in a multi-sport context. To address this gap, we introduce SportR, the first multi-sports large-scale benchmark designed to train and evaluate MLLMs on the fundamental reasoning required for sports intelligence. Our benchmark provides a dataset of 4,789 images and 2,052 videos. To enable granular evaluation, we structure our benchmark around a progressive hierarchy of question-answer (QA) pairs designed to probe reasoning at increasing depths—from simple infraction identification to complex penalty prediction. For the most advanced tasks requiring multi-step reasoning, such as determining penalties or explaining tactics, we provide 6,841 high-quality, human-authored Chain-of-Thought (CoT) annotations. In addition, our benchmark incorporates both image and video modalities and provides manual bounding box annotations to test visual grounding in the image part directly. Extensive experiments demonstrate the profound difficulty of our benchmark. State-of-the-art baseline models perform poorly on our most challenging tasks. While training on our data via Supervised Fine-Tuning and Reinforcement Learning improves these scores, they remain relatively low, highlighting a significant gap in current model capabilities. SportR presents a new challenge for the community, providing a critical resource to drive future research in multimodal sports reasoning. The dataset is available at https://github.com/chili-lab/SportR.
@@ 1 Introduction
To address these critical gaps, we introduce SportR, a new large-scale, multimodal benchmark designed to train and evaluate fine-grained reasoning for fundamental sports understanding. We focus on a diverse set of five globally popular ball and racket sports – basketball, soccer, table tennis, badminton, and American football – to provide a rich testbed for generalization. Our benchmark is based on a corpus of 4,789 images (from all five sports) and 2,052 videos (from four sports, excluding badminton), encompassing 50 distinct foul types and 12 fundamental tactics. The cornerstone of our work is a collection of 6,841 high-quality, fully human-annotated CoT. We also derive over 20,000 structured question-answer pairs.
@@ 2.2 Multimodal Sports QA Benchmarks
Question answering (QA) has become the standard way for evaluating the reasoning abilities of MLLMs across different domains Chen et al. (2025a); Liu et al. ; Yue et al. (2024); Wu et al. (2023); Shao et al. (2024a). In the sports domain, early multimodal QA benchmarks primarily focused on assessing models’ ability to recognize actions and answer descriptive or temporal questions. For instance, ActionAtlas (Salehi et al., 2024) which focuses on action recognition via multiple-choice, or Ego-Exo4D (Grauman et al., 2024) which evaluates skill proficiency and pose estimation and Sports-QA (Li et al., 2024c), which leveraged existing action recognition datasets to create a large-scale VQA benchmark, but its questions do not center on the complex, rule-based scenarios that define competitive play.
More recent benchmarks have begun to address this gap by focusing on deeper forms of reasoning. Notably, SPORTU (Xia et al., 2024b) introduced a comprehensive evaluation benchmark with a multi-level difficulty design, explicitly including “hard” questions that require understanding game rules and tactics. However, while SPORTU provides explanations for evaluation, the limitations are that there are not enough of this type of question, and these rationales were not designed as detailed and fine-grained, human-annotated reasoning processes suitable for training models to perform explicit, step-by-step reasoning. Furthermore, it lacks precise annotations for visual grounding, making it difficult to assess whether a model’s judgment is based on specific visual evidence, such as the subtle point of contact defining a foul. In addition, SPORTU primarily relies on multiple-choice questions and slow-motion replays, limiting its utility for training deep reasoning and evaluating true sports understanding. Recent studies suggest that the MCQ format may not accurately reflect an LLM’s true capabilities due to misalignment with real-world generation tasks and sensitivity to option ordering (Li et al., 2024d). A detailed comparison is provided in Appendix A.
@@ 3 SportR Benchmark
SportsImage – a dataset of 4,789 images designed to test a model’s ability to connect abstract rules to precise evidence in static scenes. Each image is paired with a CoT rationale focused on a specific foul or tactical play. This component’s progressive QA suite culminates in a unique visual grounding task, where the model is required to output the exact bounding box coordinates of the rule infraction.
SportsVideo – a dataset of 2,052 video clips created to extend the reasoning challenge to the temporal domain. This component addresses scenarios where understanding context is only possible by observing motion over time, such as a “traveling” violation in basketball. Each video is also annotated with a comprehensive CoT, from which a similar suite of progressive QA pairs. This part aims to provide a more thorough evaluation benchmark in sports understanding.
Overall, SportR contains 6,841 unique, human-authored CoT rationales, from which we derive over 20,000 structured question-answer pairs. This benchmark fills a critical gap by providing the resources to train models not just to answer what happened, but to reason about why it happened based on explicit, grounded visual evidence.
Our design prioritizes not only assessing a model’s reasoning but also verifying that this reasoning is matched to specific visual evidence. To this end, we introduce a novel explicit grounding task, where models must output the precise bounding box coordinates of a rule infraction. We focus this coordinate-based evaluation on SportsImage because fouls are often defined by discrete, localized events—such as a specific point of physical contact—that can be unambiguously annotated and evaluated in a static frame. This provides a clean and tractable setting to evaluate a skill that, as we later demonstrate, poses a significant challenge even for state-of-the-art models. Establishing a reliable baseline for this explicit form of grounding is a critical first step before tackling the greater spatio-temporal complexity of localization in video, which remains an important direction for future work. Examples of our SportsImage and SportsVideo can be found in Appendix F
@@ 3.1 Quality Control
A key principle of our data curation was to focus on fundamental and widely understood fouls and tactics, such as a basketball blocking foul or a standard post move, rather than highly specialized or controversial professional-level edge cases. This focus ensures the scenarios, while challenging for AI systems, are grounded in the established, practical expertise of our annotation team. This allows the creation of clear, consistent, and authoritative rationales, minimizing ambiguity and subjectivity.
The process began with a training phase where all annotators collaboratively developed and standardized the guidelines for writing CoT rationales and annotating grounding boxes. Specifically, annotators were trained to follow a strict “Macro-to-Micro” logical flow. The reasoning process consists of three sequential steps: (1) Identify the specific court area and involved parties; (2) Describe the action details and dynamics leading up to the event; and (3) Pinpoint the precise point of contact or critical visual evidence. Each member then worked on small batches of examples, with the group reviewing the outputs to ensure a unified standard of quality and detail before commencing the full-scale annotation.
@@ 3.2 SportsImage: Multimodal Image QA
SportsImage is the first component of our benchmark, designed to rigorously evaluate an MLLM’s ability to perform fine-grained reasoning and rule-based visual grounding in static scenes. This part covers five sports: basketball, soccer, table tennis, badminton, and American football. It comprises 4,789 images, each capturing a critical moment of gameplay across multiple sports.
Each image is paired with a comprehensive, human-authored CoT rationale. This rationale serves as the ground truth to generate a progressive QA suite of up to seven distinct questions, resulting in over 17,000 structured question-answer pairs for this component alone. These questions are structured to probe a model’s reasoning at increasing levels of depth, covering: (Q1) infraction identification, (Q2) foul classification, (Q3) penalty prediction, (Q4) free-form explanation, (Q6) offensive tactic identification, (Q7) defensive tactic identification, and culminating in (Q5) an explicit visual grounding task requiring the model to output precise bounding box coordinates.
@@ 3.3 SportsVideo: Multimodal Video QA
The dataset consists of 2,052 video clips, from which we derive over 6,000 structured question-answer pairs. Like its image-based counterpart, each video is annotated with a comprehensive, human-authored CoT rationale that explains the core foul or tactic. This CoT then serves as the ground truth for a progressive QA suite designed to evaluate spatio-temporal reasoning.
@@ 4.1.1 Model Training on SportR
Beyond zero-shot baseline evaluations, we fine-tuned an open-source model to validate SportR as an effective training resource. We employed a two-stage training process on the Qwen-2.5VL-7B model: an initial Supervised Fine-Tuning (SFT) phase followed by Reinforcement Learning (RL) using the Group Relative Policy Optimization (GRPO) algorithm (Shao et al., 2024b). The training was conducted exclusively on our SportsImage component to demonstrate its value for teaching fine-grained visual reasoning. Full details on our data preparation, SFT/RL methodology, and reward function design are deferred to Appendix B.
@@ 4.2 Evaluation
We employ two primary metrics: Intersection over Union (IoU) (Everingham et al., 2010) for the visual grounding task (Q5) and an LLM-as-Judge framework for all other text-based QA tasks (detaile prompts are in Appendix C. To mitigate the known issue of self-preference bias where models may unfairly favor their own outputs (Zheng et al., 2023), we use three proprietary models as judges: GPT-5 (a20, 2025), Gemini 2.5 Pro (Comanici et al., 2025), and Claude 4.0 Sonnet (Anthropic, 2025) and report the average score. To validate the reliability of this metric, we conducted a human verification study on a stratified subset of the test set (660 samples) using the same instruction as we provided to LLMs. We observed a strong correlation between the LLM average score and expert human judgments (Pearson $r>0.65$ across modalities), confirming that our “Average Score” approach aligns well with human evaluation. The correlation map are shown in Appendix E For each generated response, we report the scores from all three judge models as well as the average score to provide a comprehensive and robust assessment of performance.
@@ 5 Result
Our experiments are designed to demonstrate both the profound challenge posed by our benchmark and its utility as a training resource. We present results for the SportsImage and SportsVideo components separately to provide analysis of the model’s capabilities in both static and video settings.
@@ 5.1 Performance on SportsImage
The zero-shot results underscore the difficulty of our benchmark. Even the most powerful proprietary model, GPT-5, struggles to achieve high accuracy on the core reasoning tasks. Across all baselines, performance is particularly low on tasks requiring deep, specialized knowledge, such as Foul Classification (Q2) and, most notably, the Explicit Visual Grounding task (Q5), where IoU scores are consistently below 7%. This confirms that existing models lack the sports fine-grained perception and rule-based reasoning abilities that SportR is designed to measure.
@@ 5.2 Performance on SportsVideo
Table 2: Performance comparison across models (%) on SportsVideo. Q8: Video Infraction identification; Q9: Video Foul Classification; Q10: Video Penalty Prediction; Q11: Video Free-form Explanation. Q12: Offensive Tactic Identification; Q13: Defensive Tactic Identification.
@@
5.3 Error Analysis (New Section added based on reviewer feedback)
Figure 3: Error type distribution across different MLLMs on overall SPORTR tasks. The analysis reveals that Visual Perception Error is the most common issue, followed by Hallucination Error. Each error type highlights specific model limitations in comprehending the task.
To understand the specific challenges posed by SportR, we conducted a manual error analysis on 1,500 failure cases. We sampled 150 images and 150 videos from the test set for each of the six representative models. Errors were classified into five categories: (1) Visual Hallucination, fabricating non-existent objects or events; (2) Domain Knowledge Gap, misapplying sports rules or failing to recognize a standard penalty; (3) Reasoning Error, flawed logical derivation despite correct perception; (4) Format Violation, failing to follow output schema or refusing to answer; and (5) Visual Perception Error, missing critical visual evidence present in the input. The aggregate error distribution is illustrated in Figure 3. For a detailed breakdown by modality (Image vs. Video), please refer to Appendix D.
We found in video tasks, errors are overwhelmingly dominated by Visual Perception Error and Visual Hallucination. Across all models, these two categories combined frequently exceed 60-70% of total errors. This indicates that the primary barrier in video is the inability to parse fine-grained, temporal dynamics. Consequently, models rarely reach the stage of successful rule adjudication, resulting in an artificially low “Domain Knowledge Gap.”
In contrast, static image tasks—where temporal ambiguity is removed and visual perception is inherently easier—reveal the true extent of the models’ limitations in sports logic. We observe a sharp spike in Domain Knowledge Gaps when shifting from video to image. This trend is particularly evident in proprietary SOTA models. For instance, GPT-5’s Domain Knowledge Gap rises from 20.00% in video to 36.00% in images. This confirms that the decrease in visual perception errors, expected due to the static nature of images in sports, effectively unmasks the underlying deficiency in reasoning: even when SOTA models successfully perceive the visual evidence, they struggle significantly to map that evidence to the correct abstract sports rule.
@@ 6 Conclusion
In this paper, we introduced SportR, a large-scale benchmark designed to train and evaluate the fine-grained, rule-based visual reasoning of Multimodal Large Language Models. Built upon a foundation of over 6,500 fully human-authored Chain-of-Thought rationales, we aim to provide a benchmark that can be used both for training and evaluation. By integrating tasks across both images and videos, our benchmark provides a holistic assessment of a model’s ability in sports understanding. Our experiments demonstrate that while our dataset is an effective training resource, it presents a profound challenge to current models. For example, even after fine-tuning and reinforcement learning, performance on the explicit grounding task remains modest, underscoring the difficulty of connecting abstract rules to precise visual evidence. Crucially, our error analysis reveals that current MLLMs suffer from a fundamental shortage in fine-grained visual perception for dynamic events and a significant gap in aligning visual evidence with abstract domain knowledge. We hope SportR will serve as a critical tool for the community, inspiring advancements in MLLMs and contributing to more robust and reliable real-world sports understanding.
7 Acknowledgments
This project is supported in part by the Rice Ken Kennedy Institute Research Award #1081025. We thank the anonymous reviewers for their thoughtful feedback and constructive suggestions, which helped improve this work.
Appendix A Detailed Comparison with Existing Benchmarks
While benchmarks like SPORTU (Xia et al., 2024b) pioneered multi-sport evaluation, SportR represents a structural evolution designed to support training and fine-grained reasoning. The key distinctions are summarized below:
1. Training vs. Evaluation Utility.
SPORTU’s “hard” tasks rely heavily on Multiple Choice Questions (MCQs). While effective for low-cost evaluation, MCQs allow models to guess answers and cannot be used to train generative reasoning capabilities. In contrast, SportR provides over 6500 human-authored Chain-of-Thought (CoT) annotations. This transforms the benchmark from a pure evaluation set into a rich training resource that explicitly teaches models how to reason step-by-step.
2. Visual Fidelity (Normal vs. Slow Motion).
A critical limitation of SPORTU’s video subset is its reliance on slow-motion replays, which simplifies temporal perception. SportR utilizes normal-speed videos, presenting a significantly more realistic challenge. Models must capture fleeting visual cues without temporal assistance, testing true temporal understanding.
3. Task Scope: Adjudication vs. Perception.
Prior benchmarks often include basic perception tasks (e.g., counting players, identifying colors). SportR filters these out to focus exclusively on the “Reasoning Gap”: complex rule-based adjudication and visual grounding. This ensures the benchmark measures deep sports understanding rather than basic visual recognition.
Appendix B Detailed Model Training Methodology
@@
B.1 Model Training on SportR
@@
B.2 Training Data Preparation